The Garrison Company manufactures two products: Oxy Cleaner and Sonic Cleaner. The costs and revenues are as follows:

Oxy Cleaner Sonic Cleaner
Sales Price $ 80 $ 48
Variable cost per unit 30 20


Total demand for Oxy is 10,700 units and for Sonic is 6,700 units. Machine time is a scarce resource. During the year, 57,000 machine hours are available. Oxy requires 5 machine hours per unit, while Sonic requires 2.0 machine hours per unit.

What is the maximum contribution margin Garrison can achieve during a year?
Answer #1
Oxy sonic
Sales 80 48
Less: variable cost 30 20
Contribution margin per unit 50 28
Machine hour per unit 5 2
Contribution margin per hour $10 $14
Ranking 2 1

Hour required for sonic = 6700*2 = 13400 hours

Hour available for oxy = 57000-13400 = 43600 hours

Production of Oxy = 43600/5 = 8720 units

Maximum contribution margin = 8720*50 + 6700*28

= $623600